OpenStudy (ayyookyndall):

What is the equation for the line perpendicular to y= -5/3x + 11 1/3 and containing P(-2,3)? A. y-2= -3/5(x-3) B. y= -5/3x + 4 1/3 c. y= -3/5x + 4 1/5 D. y= 3/5x + 4 1/5 I think it is B..... Am I right?

OpenStudy (anonymous):

I think it's D

OpenStudy (ayyookyndall):

How so? Can you explain? @Vincentli

OpenStudy (ayyookyndall):

@amistre64 @bibby @Callisto @jim_thompson5910

OpenStudy (texaschic101):

y = -5/3x + 11 1/3 in y = mx + b form, the slope is in the m position. In this problem, the slope is -5/3. However, we are looking for a perpendicular line, so we need the negative reciprocal of the slope we found in the given equation. All that means is " flip " the slope and change the sign. So the slope we need is the negative reciprocal of -5/3 and that is 3/5 (see how I flipped the slope and changed the sign). Now we use y = mx + b slope(m) = 3/5 (-2,3)...x = -2 and y = 3 now we sub 3 = 3/5(-2) + b 3 = - 6/5 + b 3 + 6/5 = b 15/5 + 6/5 = b 21/5 = b 4 1/5 = b so your perpendicular equation is : y = 3/5x + 4 1/5...answer D

OpenStudy (anonymous):

Sorry for replying late. Was busy... But yea, @texaschic101 explained it all